# ReMA: Learning to meta-think for LLMs with multi-agent reinforcement learning

### v2

@@ 6 Conclusion
In this paper, we propose ReMA, a novel framework that elicits meta-thinking in large language models (LLMs) using multi-agent reinforcement learning. Our approach explicitly separates meta-thinking from reasoning process by assigning them to distinct high- and low-level agents. This separation enhances exploration during training and improves result interpretability. We also introduce an iterative RL training procedure tailored to our framework. We evaluate ReMA on mathematical reasoning and LLM-as-a-Judge benchmarks, testing across different base models. To ensure robustness, we design two reward mechanisms: correctness-based and consistency-based rewards. Our results demonstrate that ReMA achieves the highest performance, particularly on out-of-distribution datasets. Additionally, we provide a case study illustrating how our collaborative training strategy enables effective coordination between high- and low-level models. The success of ReMA highlights the potential of reinforcement learning and the benefits of separating metacognition from reasoning execution.
7 Limitation and Future work
This work has several limitations that open opportunities for future research:
  • Our experiments were conducted exclusively on small language models (fewer than 8B parameters), which may not fully capture the potential benefits of scaling to larger models and datasets. We plan to extend our experiments to larger language models and more extensive datasets to better evaluate scalability and performance gains.
  • The proposed MRP/MAMRP methods did not consistently outperform the VRP baseline on the tested large language models. Additionally, our formulations are currently limited to single-turn interactions. We aim to refine these methods and extend them to multi-turn settings while investigating performance variations across different model architectures.
  • Our model has not yet been evaluated on a broad range of tasks, and we have not assessed the reasoning tax, i.e. the overthinking problem in current reasoning models. A comprehensive evaluation across diverse benchmarks and tasks is necessary, including a detailed analysis of the reasoning tax to understand its impact on overall performance. And the instruction of an extra meta-thinking agent shows great potential as a regulation to prevent over-thinking.
  • The current evaluation is restricted to mathematical tasks and the LLM-as-a-judge benchmark, which may not fully represent real-world scenarios. Future studies will include systematic comparisons across multiple benchmarks to better understand relative performance and identify specific areas for improvement.
We are actively working to improve the scalability of our MARL system during both training-time RL and test-time inference, ensuring that future iterations are more robust and broadly applicable.
Appendix A Inference-time Scaling of multi-agent metacognition-integrated reasoning process
@@
Appendix B Dataset Curation of RewardBench970
[…]
JudgeBench 51.77 51.45 (-0.32) 50.65 (-1.12) 53.71 (+1.94)
Average 61.51 66.65 (+5.14) 65.53 (+4.02) 70.00 (+8.49)
JudgeBench 58.39 54.84 (-3.55) 55.81 (-2.58) 58.71 (+0.32)
Average 72.44 71.03 (-1.41) 68.06 (-4.38) 74.72 (+2.28)
Appendix C Improvement by RL Training
In this section, we illustrate the performance differences before and after RL training. We compare three inference settings:
  • VRP: A single LLM directly solves the problem step-by-step.
  • MRP: A single LLM first generates a meta-thinking message, and then solves the problem based on it.
  • MAMRP: One LLM generates the meta-thinking message, while another LLM follows it to solve the problem.
Figure 5: Relative performance gain after RL training for three different inference settings.
Figure 5 presents the results. We first evaluate these settings without RL training, i.e., using direct inference as a baseline. Then, we compare their performance with RL-trained models from Section 4.3. The results show that RL training enhances overall performance across all settings, with particularly notable improvements in MRP and MAMRP. This underscores the importance of RL fine-tuning in optimizing inference quality.
Appendix D Brief convergence analysis
We reuse the notations from Section 3.2, where $\mathbf{x}$ is task prompt, $\mathbf{y}$ is generated answer, $\mathbf{y}^{*}$ is ground-truth, $\mathbf{m}$ is metacognition on task solving, $\pi_{\theta_{h}}$ and $\pi_{\theta_{l}}$ are high- and low-level agents with parameters $\theta_{h}$ and $\theta_{l}$. We consider the joint hierarchical policy defined in Equation 8 and update the objective as in Equation 9.
$\theta_{h}^{(t+1)}=\arg\max_{\theta_{h}}J(\theta_{h},\theta_{l}^{(t)}),$
$\theta_{l}^{(t+1)}=\arg\max_{\theta_{l}}J(\theta_{h}^{(t+1)},\theta_{l}).$
Regarding the different regularizations $R_{h}$ and $R_{l}$ in Eq. 10 and 11 for the different policies, instead of directly integrating them into the loss function, we treat them as constraints, as done in Trust Region Policy Optimization (TRPO) (Schulman et al., 2015). Note that when one policy is fixed, the other policy operates in a stationary decision process.
Based on the defined objective and update method, we apply TRPO and block coordinate ascent to demonstrate that our update method theoretically converges to the optimal policies with respect to the final reward. First, recall that when updating a single policy, TRPO guarantees monotonic improvement by optimizing a lower bound. Specifically, let $\pi_{\text{old}}$ and $\pi$ represent the old and current policies, respectively. We define a surrogate objective as:
As shown by Schulman et al. (2015), the true objective of $\pi$ is lower-bounded by:
$J(\pi^{t+1})\geq J(\pi^{t}).$
$J(\theta_{h}^{t+1},\theta_{l}^{t})\geq J(\theta_{h}^{t},\theta_{l}^{t}),$
$J(\theta_{h}^{t+1},\theta_{l}^{t+1})\geq J(\theta_{h}^{t+1},\theta_{l}^{t}).$
Thus $J(\theta_{h}^{t+1},\theta_{l}^{t+1})\geq J(\theta_{h}^{t},\theta_{l}^{t})$. This repeated coordinate maximization converges to a stationary point, where no single coordinate update can further improve $J(\theta_{h},\theta_{l})$.
Given the theoretical monotonic improvement with TRPO and block coordinate ascent, we adopt a practical version of TRPO in our experiments, specifically Proximal Policy Optimization (PPO) (Schulman et al., 2017) or REINFORCE++ (Hu, 2025b).
Appendix E Detailed reward design
@@ 2. Consistency-Based Reward
These different reward formulations allow us to investigate various dimensions of metacognitive reasoning: correctness, consistency, etc. We empirically compare their effects on learned metacognitive reasoning patterns in Sec 4.4.2.
Appendix F Learning to reason from the perspective of Leader Follower Game
F.1 Leader-follower game
The leader-follower game, also known as the Stackelberg game, models interaction between two agents with parametrized strategies $\bm{w}=(\bm{w}_{1},\bm{w}_{2})$ and differentiable objective functions $(\mathcal{L}_{1},\mathcal{L}_{2}):\mathbb{R}^{d}\rightarrow\mathbb{R}$. In this framework, the leader announces its strategy first, and the follower observes this decision to respond optimally. This sequential structure enables the leader to anticipate the follower’s reaction and adjust its strategy accordingly. A Stackelberg equilibrium occurs when neither agent can unilaterally improve its objective. Denoting $\bm{w}_{1}$ as the leader’s strategy and $\bm{w}_{2}$ as the follower’s, the loss functions $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are optimized with the following bi-level structure:
$\bm{w}_{1}^{*}=\text{argmin}_{\bm{w}_{1}}\mathcal{L}_{1}(\bm{w},\bm{w}_{2}^{*}(\bm{w}_{1})),\quad\bm{w}_{2}^{*}(\bm{w}_{1})=\text{argmin}_{\bm{w}_{2}}\mathcal{L}_{2}(\bm{w}_{1},\bm{w}_{2}).$
Anil et al. (2021) apply the leader-follower game to ensure checkable answers in a prover-verifier game (PVG). The objective is a verifier that is both complete (accepts all correct proofs from a verifier) and sound (rejects all incorrect proofs from a verifier). They analyze different scenarios where the verifier acts as the leader, the prover as the follower, and both announce strategies simultaneously, forming a Nash equilibrium. The study concludes that in verifier-led PVG, a Stackelberg equilibrium is both necessary and sufficient for achieving a sound and complete verifier, whereas in other configurations, a Stackelberg equilibrium is not necessary or sufficient for this outcome.
F.2 Efficacy of LLM
Because the high-level policy possesses strong generalization capabilities, it is impractical for it to exhaustively explore every potential sub-task for each question. Instead, it naturally focuses on tasks within a feasible range of difficulty, leveraging only a limited set of coarse planning actions. Rather than pinpointing perfectly tailored sub-tasks, the policy searches for general tasks of particular computational complexity, i.e., difficulty, that it can handle reliably. Motivated by this perspective, we incorporate the concept of a reasoning boundary for large language models (LLMs) (Chen et al., 2024b). Intuitively, the reasoning boundary circumscribes the maximum difficulty of problems a model can solve at a desired accuracy level. Formally, for a model $\theta$, a task $t$, and a predefined threshold $A$, the reasoning boundary of $\theta$ represents the maximum problem difficulty $d$ that satisfies:
$\mathcal{B}_{Acc=A}(t|\theta)=\sup_{d}\{d|Acc(t|d,\theta)=A\}.$
F.3 Leader-follower Game for LLM Reasoning
To analyze the result agents by minimizing the loss functions, we adopt the completeness and soundness properties from the PVG framework for LLM reasoning. Specifically, if the high-level policy generates a sub-task sequence that is executable within the low-level policy’s capabilities, the problem must be solved (completeness). Conversely, if the sub-task sequence is incorrect or beyond the low-level policy’s capacity, the problem cannot be solved (soundness). To achieve this, we utilize the conclusion from Anil et al. (2021), which positions the low-level policy as the leader and the high-level policy as the follower, equilibria guarantee the complete and sound low-level policy.
where we can compute the difficulty $d_{k}$ from $t_{k}$ and $x$. where the difficulty $d_{k}$ can be derived from $t_{k}$ and $x$. The loss function in Equation 13 ensures that the selected sub-tasks are optimal for the low-level policy. Here we provide a theoretical condition under which the most efficient solution strategy can be identified, according to the efficacy of the LLM.
By framing the problem-solving process as a leader-follower game, with the low-level policy designated as the leader, we can construct a bi-level optimization problem to identify an equilibrium. Following the formulation in Section F.1, the problem is expressed as:
Then we can apply bi-level optimization techniques.
Appendix G Pseudocode of ReMA
@@
Appendix H Supervised fine-tuning data collection

### v3

@@ 5 Conclusion
In this paper, we introduced ReMA, a novel framework that leverages multi-agent reinforcement learning to elicit meta-thinking in large language models. By explicitly separating meta-thinking and reasoning processes into distinct agents, our approach enhances both exploration during training and the interpretability of model outputs. We tailored RL algorithms and reward functions to ensure reliable performance. Through comprehensive experiments on mathematical reasoning and LLM-as-a-Judge benchmarks, ReMA consistently achieved superior results, particularly on out-of-distribution datasets. We further extend ReMA to multi-turn settings, enabling the framework to handle more complex reasoning scenarios that require more communication between agents. Our ablations demonstrate how effective coordination between agents evolves, highlighting the promise of reinforcement learning and structured agents’ collaboration for advancing the capabilities of language models in complex reasoning tasks.
Appendix Table of Contents
  • A Related work A
    • A.1 Single LLM Reasoning A.1
    • A.2 Multiple LLM Reasoning A.2
    • A.3 Hierarchical Reasoning A.3
    • A.4 RL in LLM A.4
  • B Limitation and Future Work B
  • C Supplementary Materials for Method in Section 3 C
    • C.1 Inference-time Scaling For ReMAC.1
    • C.2 Detailed reward design C.2
    • C.3 Pseudocode of ReMAC.3
    • C.4 Brief convergence analysis C.4
    • C.5 Learning to reason from the perspective of Leader Follower Game C.5
  • D Training Details D
  • E Other Experiments E
    • E.1 Reward functions shape cross-agent behaviors E.1
    • E.2 Detailed Training Curves on Different Datasets of Multi-turn ReMAE.2
  • F Qualitative results F
    • F.1 High-level policy finds better plans F.1
    • F.2 Case study for Experiments of Different Reward Functions in Appendix E.1 F.2
    • F.3 Case study for Adaptive Meta-thinking in Single-Turn ReMA in Section 4.2.2 F.3
  • G Prompts G
Appendix A Related work
Drawing from the bitter lesson [Sutton, 2019], two methods that appear to scale effectively are searching and learning, aligning with current trends in large language models [Xu et al., 2025]. At present, researchers are leveraging these methods to maximize the capabilities of individual transformers, while other efforts are exploring architectures that involve multiple interacting entities. In this paper, we examine this divergence within the context of LLM reasoning, a capability that allows large language models to solve problems through logical reasoning, step-by-step analysis, and inference [Wang et al., 2024a].
A.1 Single LLM Reasoning
Main research works in reasoning involving a single LLM utilize search-based and post-training methods. The fundamental elements of searching methods are text generation and evaluation. Generation schemes include In-Context Learning [Brown et al., 2020], Beam Search [Graves, 2012], and various tree-based searching [Snell et al., 2024]; Evaluation approaches often use outcome accuracy, self-consistency [Wang et al., 2022], or process reward signal [Lightman et al., 2023] as the criteria to select high-quality responses from the generated texts. Post-training method is another research line in opposition to pre-training. Popular training pipelines often involve specific data construction followed by Supervised Fine-tuning [Qin et al., 2024, Ouyang et al., 2022, Hui et al., 2024, Liu et al., 2024], or reinforcement learning to interactively explore learning patterns [Wang et al., 2024a, Zhang et al., 2024a, DeepSeek-AI et al., 2025, Xu et al., 2025].
A.2 Multiple LLM Reasoning
Integrating multiple entities can potentially surpass the intelligence of the individual model [Chen et al., 2023]. With the rapid emergence of large language models showing a varying level of abilities, some studies have explored facilitating discussions among multiple off-the-shelf LLMs [Zhang et al., 2025a, Chen et al., 2024a, Wang et al., 2023, Du et al., 2023, Zhuge et al., 2023, Tang et al., 2023, Hao et al., 2025, Akata et al., 2023, Hong et al., 2023, Zhang et al., 2024b], taking the form of free discussion [Du et al., 2023, Liang et al., 2023] or structured role assignments [Hong et al., 2023, Zhang et al., 2024b]. Some have applied routing mechanisms to assign tasks to the most suitable expert models [Hu et al., 2024b, Stripelis et al., 2024, Ding et al., 2024, Yue et al., 2025a, Chen et al., 2024c] or merging mechanisms to develop more versatile models [Yadav et al., 2024, Yu et al., 2024, Zhang et al., 2025b]. Beyond aggregating static knowledge from multiple agents, multi-agent LLM training can also enhance reasoning capabilities. For example, multi-agent debates can generate diverse synthetic data, which can subsequently be used for supervised fine-tuning [Estornell et al., 2024, Li et al., 2024, Motwani et al., 2024, Dong and Ma, 2025, Perez et al., 2022, Ye et al., 2025a, Subramaniam et al., 2025]. Reinforcement learning (RL) methods have also been adopted to improve LLM reasoning in areas such as alignment [Perez et al., 2022, Ma et al., 2023] and legibility [Kirchner et al., 2024]. Motwani et al. [2024] utilize a three-agent system for generation and fine-tune the models using Direct Preference Optimization (DPO). Reinforcement Learning with Generative Reward Models (GenRM) [Mahan et al., 2024, Ye et al., 2025b, Jiao et al., 2024, Wang et al., 2024b] represents another common approach of multi-agent training, where the reward signal is derived from the token probabilities of another LLM, coupled with the reasoning process. While our work aligns with these efforts, it diverges by using an additional tunable LLM to provide metacognitive instructions, guiding the low-level LLM during learning, rather than relying on a static GenRM. The most closely related works to ours are MAPoRL [Park et al., 2025] and COPRY [Ma et al., 2024]. MAPoRL is a multi-agent debating framework that uses multi-agent reinforcement learning (MARL) with a learned verifier to fine-tune each LLM agent. COPRY duplicates an LLM into two agents, training them simultaneously in the roles of pioneer and observer using RL. Shen et al. [2025] trained with a novel Chain-of-Action-Thought (COAT) framework that embeds meta-action tokens for self-reflection and exploration into an autoregressive search process. However, unlike our approach, which explicitly separates metacognition from plan execution, these methods do not decompose the reasoning process but instead focus on improving direct chain-of-thought generation. Furthermore, our experiments are conducted on a larger scale and include more challenging problems.
A.3 Hierarchical Reasoning
Partitioning reasoning into hierarchical processes has been explored in prior research to make biological sense [Ye et al., 2018, Langley et al., 2004]. In the context of language models, a hierarchical structure has been used to facilitate diverse reasoning patterns, including planning [Puerta-Merino et al., 2025, Sun et al., 2024, Song et al., 2023, Rana et al., 2023, Chen et al., 2024d, Yan et al., 2023, Xiao et al., 2024], validation [Haji et al., 2024, Xi et al., 2024] and self-refinement [Madaan et al., 2023, Kumar et al., 2024, Welleck et al., 2022]. For instance, EvalPlanner [Saha et al., 2025b] is a framework that conducts reasoning through plan generation and execution. DOTS [Yue et al., 2024] extends decomposition by integrating a tree-based searching method with Analysis, Solution, and Verification layers. Marco-o1 [Zhao et al., 2024] focuses on open-ended problem-solving and abstract thinking, dynamically adjusting reasoning granularity and incorporating reflection mechanisms to enhance reasoning performance. Beyond these approaches, metacognition [Flavell, 1979] has been identified as another critical component of reasoning, referring to the intuitive understanding of one’s own cognitive and reasoning processes [Gao et al., 2024, Wang and Zhao, 2023]. Wang and Zhao [2023] proposed a metacognitive prompting strategy to improve large language model (LLM) capabilities. Didolkar et al. [2024] further developed a prompt-guided method that enables models to label math problems with the required skills and subsequently use these labels to solve new problems. Gao et al. [2024] introduce meta-reasoner which use contextual multi-arm bandit to learn a high-level “advisor” over low-level reasoning process. Xiang et al. [2025] provides a Meta-CoT framework to think about its own thinking. They use construction-based methods as well as reinforcement learning to develop meta-cognitive skills. Qingsong et al. [2025] introduces a RL framework for dynamic instruction selection during fine-tuning. In our work, we also value reflecting on reasoning processes, and we enhance metacognitive abilities through two-agent interaction and reinforcement learning at both end.
A.4 RL in LLM
Recent advancements in applying RL to LLMs have enhanced their reasoning and decision-making capabilities. Liu et al. [2025] examines token-level optimization biases by introducing Dr. GRPO to stabilize policy gradients. VAPO [Yue et al., 2025b] enhances PPO with value-aware perturbations and adaptive reward shaping to improve robustness in sparse-reward reasoning tasks. DAPO [Yu et al., 2025] provides a scalable, modular RL framework that integrates distributed rollout collection and dynamic replay buffers for reproducible training at scale. SimpleRL-Zoo [Zeng et al., 2025] conducts zero-shot RL experiments across open-base LLMs to uncover emergent cognitive behaviors under minimal reward signals. Echo Chamber [Zhao et al., 2025] investigates how RL fine-tuning algorithms can amplify pretrained model biases and proposes regularization to mitigate over-amplification. Wen et al. [2024] decomposes high-level language actions into token-level operations to achieve finer-grained credit assignment. Some works push RL training for single-turn to multi-turn. Search-R1 [Jin et al., 2025] trains LLMs to orchestrate multi-turn search strategies with RL-optimized decision policies to improve question-answering accuracy. ArCHer [Zhou et al., 2024] employs a hierarchical, multi-turn RL architecture with manager and worker policies to efficiently handle long-horizon dialogue tasks. RAGEN [Wang et al., 2025] introduces trajectory filtering and critic modules within a multi-turn RL framework to stabilize learning and reduce shallow policy behaviors.
Appendix B Limitation and Future Work
In this work, we only test ReMA on math and LLM-as-a-Judge benchmarks. Though the results shows the effectiveness of ReMA, adopting ReMA to tasks where naturally needs multi-turn interaction between several interleaved agents has great potential. Moreover, a comprehensive understanding of the learning dynamics of multi-turn RL and multi-turn MARL for LLMs is needed. Finally, there’s still sufficient space to further improve the procedure of multi-turn multi-agent rollout through modern LLM speed up techniques, e.g. prefill-decode disaggregation and asynchronous rollout.
Appendix C Supplementary Materials for Method in Section 3
C.1 Inference-time Scaling of ReMA
@@
C.2 Detailed reward design
As described in Sec. 3.2, we update both high-level and low-level agents by assigning rewards based on the low-level policy output. Below, we outline several potential reward designs:
@@ 2. Consistency-Based Reward
These different reward formulations allow us to investigate various dimensions of metacognitive reasoning: correctness, consistency, etc. We empirically compare their effects on learned metacognitive reasoning patterns in Sec. E.1.
C.3 Pseudocode of ReMA
@@
C.4 Brief convergence analysis
We reuse the notations from Sec. 3.2, where $\mathbf{x}$ is task prompt, $\mathbf{y}$ is generated answer, $\mathbf{y}^{*}$ is ground-truth, $\mathbf{m}$ is metacognition on task solving, $\pi_{\theta_{h}}$ and $\pi_{\theta_{l}}$ are high- and low-level agents with parameters $\theta_{h}$ and $\theta_{l}$. We consider the joint hierarchical policy defined in Eq. (8) and update the objective as in Eq. (9).
$\theta_{h}^{(t+1)}=\arg\max_{\theta_{h}}J(\theta_{h},\theta_{l}^{(t)}),$
$\theta_{l}^{(t+1)}=\arg\max_{\theta_{l}}J(\theta_{h}^{(t+1)},\theta_{l}).$
Regarding the different regularizations $R_{h}$ and $R_{l}$ in Eqs. (10) and (11) for the different policies, instead of directly integrating them into the loss function, we treat them as constraints, as done in Trust Region Policy Optimization (TRPO) [Schulman et al., 2015]. Note that when one policy is fixed, the other policy operates in a stationary decision process.
Based on the defined objective and update method, we apply TRPO and block coordinate ascent. First, recall that when updating a single policy, TRPO guarantees monotonic improvement by optimizing a lower bound. Specifically, let $\pi_{\text{old}}$ and $\pi$ represent the old and current policies, respectively. We define a surrogate objective as:
As shown by Schulman et al. [2015], the true objective of $\pi$ is lower-bounded by:
$J(\pi^{t+1})\geq J(\pi^{t}).$
$J(\theta_{h}^{t+1},\theta_{l}^{t})\geq J(\theta_{h}^{t},\theta_{l}^{t}),$
$J(\theta_{h}^{t+1},\theta_{l}^{t+1})\geq J(\theta_{h}^{t+1},\theta_{l}^{t}).$
Thus $J(\theta_{h}^{t+1},\theta_{l}^{t+1})\geq J(\theta_{h}^{t},\theta_{l}^{t})$. This repeated coordinate maximization converges to a fixed point, where no single coordinate update can further improve $J(\theta_{h},\theta_{l})$.
Given the theoretical monotonic improvement with TRPO and block coordinate ascent, we adopt a practical version of TRPO in our experiments, specifically Proximal Policy Optimization (PPO) [Schulman et al., 2017] or GRPO [Shao et al., 2024].
C.5 Learning to reason from the perspective of Leader Follower Game
C.5.1 Leader-follower game
The leader-follower game, also known as the Stackelberg game, models interaction between two agents with parametrized strategies $\bm{\theta}=(\bm{\theta}_{1},\bm{\theta}_{2})$ and differentiable objective functions $(\mathcal{L}_{1},\mathcal{L}_{2}):\mathbb{R}^{d}\rightarrow\mathbb{R}$. In this framework, the leader announces its strategy first, and the follower observes this decision to respond optimally. This sequential structure enables the leader to anticipate the follower’s reaction and adjust its strategy accordingly. A Stackelberg equilibrium occurs when neither agent can unilaterally improve its objective. Denoting $\bm{\theta}_{1}$ as the leader’s strategy and $\bm{\theta}_{2}$ as the follower’s, the loss functions $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are optimized with the following bi-level structure:
$\bm{\theta}_{1}^{*}=\text{argmin}_{\bm{\theta}_{1}}\mathcal{L}_{1}(\bm{\theta},\bm{\theta}_{2}^{*}(\bm{\theta}_{1})),\quad\bm{w}_{2}^{*}(\bm{\theta}_{1})=\text{argmin}_{\bm{\theta}_{2}}\mathcal{L}_{2}(\bm{\theta}_{1},\bm{\theta}_{2}).$
Anil et al. [2021] apply the leader-follower game to ensure checkable answers in a prover-verifier game (PVG). The objective is a verifier that is both complete (accepts all correct proofs from a verifier) and sound (rejects all incorrect proofs from a verifier). They analyze different scenarios where the verifier acts as the leader, the prover as the follower, and both announce strategies simultaneously, forming a Nash equilibrium. The study concludes that in verifier-led PVG, a Stackelberg equilibrium is both necessary and sufficient for achieving a sound and complete verifier, whereas in other configurations, a Stackelberg equilibrium is not necessary or sufficient for this outcome.
C.5.2 Efficacy of LLM
Because the high-level policy possesses strong generalization capabilities, it is impractical for it to exhaustively explore every potential sub-task for each question. Instead, it naturally focuses on tasks within a feasible range of difficulty, leveraging only a limited set of coarse planning actions. Rather than pinpointing perfectly tailored sub-tasks, the policy searches for general tasks of particular computational complexity, i.e., difficulty, that it can handle reliably. Motivated by this perspective, we incorporate the concept of a reasoning boundary for large language models (LLMs) [Chen et al., 2024b]. Intuitively, the reasoning boundary circumscribes the maximum difficulty of problems a model can solve at a desired accuracy level. Formally, for a model $\theta$, a task $t$, and a predefined threshold $A$, the reasoning boundary of $\theta$ represents the maximum problem difficulty $d$ that satisfies:
$\mathcal{B}_{Acc=A}(t|\theta)=\sup_{d}\{d|Acc(t|d,\theta)=A\}.$
C.5.3 Leader-follower Game for LLM Reasoning
To analyze the result agents by minimizing the loss functions, we adopt the completeness and soundness properties from the PVG framework for LLM reasoning. Specifically, if the high-level policy generates a sub-task sequence that is executable within the low-level policy’s capabilities, the problem must be solved (completeness). Conversely, if the sub-task sequence is incorrect or beyond the low-level policy’s capacity, the problem cannot be solved (soundness). To achieve this, we utilize the conclusion from Anil et al. [2021], which positions the low-level policy as the leader and the high-level policy as the follower, equilibria guarantee the complete and sound low-level policy.
where we can compute the difficulty $d_{k}$ from $t_{k}$ and $x$. where the difficulty $d_{k}$ can be derived from $t_{k}$ and $x$. The loss function in Eq. (15) ensures that the selected sub-tasks are optimal for the low-level policy. Here we provide a theoretical condition under which the most efficient solution strategy can be identified, according to the efficacy of the LLM.
By framing the problem-solving process as a leader-follower game, with the low-level policy designated as the leader, we can construct a bi-level optimization problem to identify an equilibrium. Following the formulation in Sec. C.5.1, the problem is expressed as:
Then we can apply bi-level optimization techniques.
Appendix D Training Details
D.1 Single-turn ReMA
We refer to Appendix G for prompts we use during training. We implement the training pipeline with OpenRLHF [Hu et al., 2024a] which is a highly efficient codebase and is easy to scale up. We select REINFORCE++ to save resources and for efficient training. All experiments are conducted in a node of 8 NVIDIA A100 GPUs. We use bf16, Zero2, Flash-Attention and gradient checkpointing to run our experiments.
D.1.1 Supervised fine-tuning data collection